# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v2

@@ 3.4 Stable & MoE Expanded Framework
where $l_{\text{MoE}}$ represents the layer starts to utilize MoE. The converted $\phi^{\prime}=\phi+\Delta\phi_{\text{LoRA}}^{\text{(MHA)}}$, $\gamma^{\prime}=\gamma+\Delta\gamma_{\text{LoRA}}^{\text{(FFN)}}$. For MoE layer, we initialize $\mathcal{N}+1$ experts with the weight of the pre-trained FFN $\gamma$. Here, it concludes $\mathcal{N}$ reconcile experts to learn specialized knowledge and $1$ anchor experts to learn the common knowledge. Let $E_{\gamma_{i}}$ denotes the $i$-th expert, $i=1\dots\mathcal{N}+1$. And $\gamma_{i}$ is the parameter of the $i$-th expert, at the beginning of the training, these experts have identical weights, i.e., $\gamma_{1}=\gamma_{2}=\cdots=\gamma$. Router $R_{\psi}$ is random initialized with Kaiming uniform (He et al., 2015), where $\psi$ is the learnable parameter of the router.
Let $s$ denote the output of the router logits for an input token, assume $\mathcal{E}$ experts are chosen, the output of the MoE layer can be written as:
$\mathbf{x_{t}^{\prime}}=\sum_{i=1}^{\mathcal{N}}\text{TopK}(s,\mathcal{E})E_{\phi_{i}}(\mathbf{x_{t}})+E_{\phi_{\mathcal{N}+1}}(\mathbf{x_{t}})$ (9)
3.5 Optimization
Training strategy of Omni-Mol consists of two stages.
Stage 1: We perform multimodal alignment on PubChem (Kim et al., 2022), learning to describe molecules through graph modality features. The input consists of instructions and graph data, excluding SELFIES. Only the multimodal projector $f_{p}$ is trainable.
Stage 2: We fine-tune Omni-Mol by freezing the pre-trained parameters (wrapped by PEFT adapters), while the adapters, the MoE layers and the multimodal projector stay active throughout fine-tuning.
@@ 3.5 Optimization
For stage 2, we incorporate an additional auxiliary load balancing loss for the MoE layers, assume an input tensor $x\in\mathbb{R}^{B\times T\times d}$, and $\mathcal{E}$ experts out of $\mathcal{N}$ is selected, the load balancing loss is: $\mathcal{L}_{\text{aux}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{\mathcal{N}}C_{ij}\cdot\bar{s}_{ij}$, where $C_{ij}=\frac{\mathcal{N}}{T\mathcal{E}}\sum_{t=1}^{T\mathcal{E}}\mathbf{1}\{\text{t'th token selects expert j}\},\bar{s}_{ij}=\frac{1}{T}\sum_{t=1}^{T}s_{i,j,t}$ and $\mathbf{1}\{\cdot\}$ is an indicator function. This load balancing loss used in Liu et al. (2024a) additionally considers the sequence-level information.
@@
3.6 Theoretical Analysis
In theory, the regularization effect of omni-molecular tasks training can enable Omni-Mol to learn more general representations. However, demonstrating this phenomenon empirically remains challenging. Our key insight is that as more tasks are learned together, the solution space of the problems becomes progressively smaller. Assume the hypothesis space $\mathcal{V}$ of the model, and the solution of a task $i$ is $\mathcal{F}_{i}$, where $\mathcal{F}_{i}\subseteq\mathcal{V}$, for omni-molecular tasks learning with $n$ different tasks, the solution space will be: $\mathcal{F}_{\text{general}}^{(n)}=\bigcap_{i=1}^{n}\mathcal{F}_{i}$.
Theorem 3.1.
For $m>n,m,n\in\mathbb{Z}^{+}$, we have $\mathcal{F}_{\text{general}}^{(m)}\subseteq\mathcal{F}_{\text{general}}^{(n)}$. If $\forall n\leq i,j\leq m,i\neq j$, we have $\mathcal{F}_{i}\neq\mathcal{F}_{j}$, then $\mathcal{F}_{\text{general}}^{(m)}\subsetneq\mathcal{F}_{\text{general}}^{(n)}$.
The proof of Theorem 3.1 can be found in Appendix H.1
Therefore, as the number of tasks $n$ increases, the similarity between the representations learned from the solution space of $n-1$ tasks and the solution space of $n$ tasks will become increasingly higher. Let $R_{n}\in\mathcal{F}_{\text{general}}^{(n)}$ denotes the representation learned from a solution space, consider a series of learned representation $S_{R}=\{R_{1},R_{2},\dots,R_{n}\}$ learned from $S_{\mathcal{F}}=\{F_{\text{general}}^{(1)},F_{\text{general}}^{(2)},\dots,F_{\text{general}}^{(n)}\}$. Assume function $\Gamma(\cdot,\cdot)$ measures the similarity between two representations. we expect that, $\exists N\in\mathbb{Z}^{+}$, such that $\forall n>N$, $\exists i<N$, we have $\Gamma(R_{n},R_{n+1})>\Gamma(R_{i},R_{i+1})$.
In our experimental validation, we construct a series of mixed datasets comprising 1, 2, 4, and 8 tasks to form the solution space sequence $S_{\mathcal{F}}$. We train the model on each of these datasets and subsequently extract representations for a specific task. Finally, we calculate the similarity between representations $R_{i}$ and $R_{j}$ for $i,j=0,\dots,3$. See Appendix C.5 for the calculation of similarity.
Model Type #Param Exact $\uparrow$ BLEU $\uparrow$ Levenshtein $\downarrow$ RDK $\uparrow$ MACCS $\uparrow$ Morgan $\uparrow$ Validity $\uparrow$
@@ 3.6 Theoretical Analysis
Vicuna (Zheng et al., 2023b) In-Context Learning 6.7B 0.000 0.057 41.690 0.007 0.016 0.006 0.059
LLaMA2 (Touvron et al., 2023) Specialist(PEFT) 6.7B 0.012 0.804 29.947 0.499 0.649 0.407 1.000
Mol-Instruction (Fang et al., 2024) Specialist(PEFT) 6.7B 0.045 0.654 27.262 0.313 0.509 0.262 1.000
HIGHT (Chen et al., 2024) Specialist(PEFT) 6.7B 0.293 0.935 16.687 0.774 0.618 0.566 1.000
InstructMol (Cao et al., 2025) Specialist(PEFT) 6.7B 0.536 0.967 10.851 0.776 0.878 0.741 1.000
PRESTO (Cao et al., 2024) Generalist 3.2B 0.691 0.976 6.525 0.871 0.931 0.841 1.000
Omni-Mol Generalist 1.7B 0.718 0.981 6.528 0.878 0.934 0.854 1.000
@@
Vicuna (Zheng et al., 2023b) In-Context Learning 6.7B 0.000 0.057 46.877 0.025 0.030 0.021 0.017
LLaMA2 (Touvron et al., 2023) Specialist(PEFT) 6.7B 0.000 0.283 53.510 0.136 0.294 0.106 1.000
Mol-Instruction (Fang et al., 2024) Specialist(PEFT) 6.7B 0.009 0.705 31.227 0.283 0.487 0.230 1.000
HIGHT (Chen et al., 2024) Specialist(PEFT) 6.7B 0.202 0.914 20.194 0.772 0.623 0.577 0.999
InstructMol (Cao et al., 2025) Specialist(PEFT) 6.7B 0.407 0.941 13.967 0.753 0.852 0.714 1.000
PRESTO (Cao et al., 2024) Generalist 3.2B 0.531 0.958 10.298 0.823 0.887 0.790 1.000
Omni-Mol Generalist 1.7B 0.559 0.961 9.263 0.840 0.900 0.809 1.000
@@
Vicuna (Zheng et al., 2023b) In-Context Learning 6.7B 0.000 0.010 27.948 0.038 0.002 0.001 0.007
LLaMA2 (Touvron et al., 2023) Specialist(PEFT) 6.7B 0.000 0.283 53.510 0.136 0.294 0.106 1.000
Mol-Instruction (Fang et al., 2024) Specialist(PEFT) 6.7B 0.044 0.224 23.167 0.237 0.364 0.213 1.000
HIGHT (Chen et al., 2024) Specialist(PEFT) 6.7B 0.067 0.482 27.167 0.462 0.346 0.303 1.000
InstructMol (Cao et al., 2025) Specialist(PEFT) 6.7B 0.129 0.610 19.664 0.444 0.539 0.400 1.000
PRESTO (Cao et al., 2024) Generalist 3.2B 0.212 0.712 16.313 0.544 0.607 0.479 1.000
Omni-Mol Generalist 1.7B 0.257 0.763 13.558 0.601 0.660 0.556 1.000
Table 1: Comprehensive comparisons on three reaction tasks. PEFT is short for parameter-efficient fine-tuning. PRESTO represents our re-implementation based on source codes.
4 Experiments
We aim to address the following concerns: (1) Compared with existing baselines, can Omni-Mol achieve the best performances on the comprehensive omni-molecular datasets with 15 tasks simultaneously? (2) Is Omni-Mol a scalable framework with the capacity and potential to solve complex molecular tasks? (3) Are all key components of Omni-Mol essential for solving conflict collapse? (4) How can we verify that Omni-Mol converges reliably and progressively refines its representations toward a universal convergent space? We begin by describing the experimental setup, then answer all the questions in the subsequent sections.
4.1 Setup and Baselines
Datasets. To construct a general-purpose model, we select 15 tasks across 4 categories to cover as many diverse tasks as possible. Here, we briefly introduce the datasets.
  • Reaction Tasks: Forward Prediction ($\sim$124k), Retrosynthesis ($\sim$128k), Reagent Prediction ($\sim$124k), Solvent Prediction ($\sim$67k), Catalyst Prediction ($\sim$10k).
  • Regression Tasks: LogP Prediction ($\sim$10k), SCF Prediction ($\sim$623k), TPSA Prediction ($\sim$11k), Complecity Prediction ($\sim$11k), Molecular Weight Prediction ($\sim$11k), Yield Regression ($\sim$9k) from, Quantum Mechanics Property Prediction ($\sim$36k).
  • Description Tasks: Molecular Caption (Molcap) ($\sim$29k) and Description QA ($\sim$56k).
  • Action Tasks: Experiment Procedure Prediction ($\sim$241k).
Details of datasets can be found in Appendix B.
Baselines. To ensure a fair comparison, we first choose representative LLM-based models such as InstructMol and HIGHT, and also report several previous baselines, including Mol-Instruction, Llama, Vicuna, among others, some of which are derived through In-Context Learning. For datasets with fewer models, we re-implement PRESTO as baseline.
@@ 4.1 Setup and Baselines
GPT-3.5-turbo1 Retrieval - 0.565 0.482 0.623 0.450 0.543 0.585
GPT-4-03141 Retrieval - 0.607 0.525 0.634 0.476 0.562 0.610
BioMedGPT2 Generalist 10B 0.234 0.141 0.386 0.206 0.332 0.308
Mol-Instruction Specialist 6.7B 0.249 0.171 0.331 0.203 0.289 0.271
HIGHT Specialist 6.7B 0.498 0.397 0.582 0.414 0.518 0.525
InstructMol Specialist 6.7B 0.475 0.371 0.566 0.394 0.502 0.509
Omni-Mol Generalist 1.7B 0.544 0.456 0.610 0.456 0.549 0.579
@@
Llama23 Specialist 6.7B 0.282 0.232 0.351 0.221 0.304 0.469
3D-MoLM(S)4 Specialist 6.7B 0.320 0.261 0.401 0.256 0.346 0.522
3D-MoLM(G)4 Generalist 6.7B 0.318 0.261 0.401 0.259 0.350 0.519
Omni-Mol Generalist 1.7B 0.516 0.440 0.529 0.382 0.492 0.580
Table 2: Main results of molecular captioning and description QA task, 1: 10-shot results from Li et al. (2024a), 2,3: results from Luo et al. (2023); Li et al. (2024b), 4: (S,G) means the specialist and generalist version of 3D-MoLM separately. B: BLEU, R: ROUGE, M: METEOR.

### v3

@@ 5.1 Gradient Adaptive LoRA
We conduct experiments on tasks across different domains in an attempt to identify an appropriate optimal rank. As shown in Figure 3, we observe that the performance of each domain-specific task varies under different rank settings, and the optimal rank also differs across domains. For instance, the optimal rank for forward prediction is 128, whereas for molcap it is 32. We attribute this to the differences in the intrinsic dimensionality of attention weights across tasks aghajanyan2020intrinsic . Therefore, employing a static adapter to simultaneously learn multiple molecular tasks may be suboptimal.
During multi‑task training, certain tasks provide highly informative signals, whereas others contribute predominantly redundant information. Consequently, it is desirable to introduce a dynamic coefficient that adaptively amplifies the gradients associated with each task throughout the training process.
We propose an adaptive adapter, Gradient Adaptive LoRA (GAL), which introduces a dynamic scaling factor to modulate the fusion of the updated weights, $\gamma_{\theta}=\alpha/r^{p}+\beta$, where $\theta=\{\alpha,p,\beta\}$ are learnable parameters. Here, the $p$ exponent lets us model rank effects and $\beta$ can provide a direct adjustment to the scaling factor. This simple yet effective modification enables the adapter to dynamically adjust its scaling factor during training, allowing it to better adapt to the intrinsic dimension of the data.
Through this approach, we can dynamically adjust the amplitude of the gradient. The backward propagation of a LoRA on the downstream fine-tuning data $\mathcal{D}$ is modified as:
@@
Implementation details are provided in Appendix C.
5.2 Mixture-of-GAL-Experts (MoGE) Expansion
Omni-Mol needs to learn a wide range of different tasks and handle multiple modalities, including graph features, text, and SELFIES. While SELFIES is treated as regular text input to the LLM, it inherently differs significantly from natural language semantics, requiring the model to separately learn how to understand and generate SELFIES expressions.
Figure 3: Evidence of varying intrinsic dimensions across task-specific representations. We observe that the optimal LoRA rank (indicated by the yellow triangle) differs across tasks.
We aim for the model to simultaneously learn general knowledge while also differentiating different modalities and tasks. Hence, we propose Mixture-of-GAL-Experts (MoGE). We borrow the idea of MoE dai2024deepseekmoe and perform upcycling komatsuzaki2023sparse ; lin2024moe with the aforementioned GAL. We first construct $\mathcal{N}$ routed experts, each targeting specialized knowledge areas, and dynamically balance conflicting signals among these experts to effectively mitigate task-level conflicts. Besides, we introduce an additional shared expert to learn the common knowledge that underpins fundamental understanding across tasks, by consistently capturing and aligning shared features to maintain a stable global representation.
Specifically, we modify the Multi-Head Attention(MHA) and Feed Forward Network(FFN) layer as:
[…]
where $h_{l}$ is the hidden states of the l’th layer, $\phi$, $\rho$ are the pre-trained parameters of the LLM, LN represents the norm layer of the LLM. We wrap the parameter $\rho^{\prime}=\rho+\text{GAL}_{\rho}$, $\phi^{\prime}=\phi+\text{GAL}_{\phi}$. For MoGE layer, we initialize $\mathcal{N}+1$ experts with the weight of the pre-trained FFN $\rho$. Here, it concludes $\mathcal{N}$ routed experts to learn specialized knowledge and $1$ shared experts to learn the common knowledge. Let $\rho_{i}$ be the parameter of the $i$-th expert, and at the beginning of the training, these experts have identical weights, i.e., $\rho_{1}=\rho_{2}=\cdots=\rho$. Router $R_{\psi}$ is randomly initialized with Kaiming uniform he2015delving , where $\psi$ is the learnable parameter of the router.
5.3 Training
The training strategy of Omni-Mol consists of two stages.
Stage 1: We perform multimodal alignment on PubChem PubChem , learning to describe molecules through graph modality features. The input consists of instructions and graph data, excluding SELFIES. Only the multimodal projector $f_{p}$ is trainable.
Stage 2: We fine-tune Omni-Mol by freezing the pre-trained parameters (wrapped by GAL), while the adapters, expert router, and the multimodal projector stay active throughout fine-tuning.
@@ 5.3 Training
For stage 2, we incorporate an additional auxiliary load balancing loss for the MoGE layers, assume an input tensor $x\in\mathbb{R}^{B\times T\times d}$, and $\mathcal{E}$ experts out of $\mathcal{N}$ is selected, the load balancing loss is: $\mathcal{L}_{\text{aux}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{\mathcal{N}}C_{ij}\cdot\bar{s}_{ij}$, where $C_{ij}=\frac{\mathcal{N}}{T\mathcal{E}}\sum_{t=1}^{T\mathcal{E}}\mathbf{1}\{\text{t'th token selects expert j}\},\bar{s}_{ij}=\frac{1}{T}\sum_{t=1}^{T}s_{i,j,t}$ and $\mathbf{1}\{\cdot\}$ is an indicator function. Here, $s_{i,j,t}$ is the router logit of the t’th token for j’th expert in batch i. This load balancing loss used in liu2024deepseek additionally considers the sequence-level information.
@@
6 Experiments
Model #Par Exa BLEU Lev RDK MAC Mor Val
@@ 6 Experiments
DeepSeekV3 zheng2023judging ICL 685B 0.35 0.939 12.76 0.719 0.823 0.68 1.00
Llama2 touvron2023llama SL 6.7B 0.01 0.804 29.95 0.499 0.649 0.41 1.00
Mol-Ins fang2024molinstructions SL 6.7B 0.05 0.654 27.26 0.313 0.509 0.26 1.00
HIGHT chen2024hight SL 6.7B 0.29 0.935 16.69 0.774 0.618 0.57 1.00
InstructMol cao2023instructmol SL 6.7B 0.54 0.967 10.85 0.776 0.878 0.74 1.00
PRESTO cao-etal-2024-presto GL 3.2B 0.69 0.976 6.53 0.871 0.931 0.84 1.00
Omni-Mol GL 2.2B 0.73 0.980 5.55 0.895 0.947 0.87 1.00
Model #Par Exa BLEU Lev RDK MAC Mor Val
@@
DeepSeekV3 zheng2023judging ICL 685B 0.29 0.93 14.32 0.725 0.827 0.68 1.00
Llama2 touvron2023llama SL 6.7B 0.00 0.283 53.51 0.136 0.294 0.11 1.00
Mol-Ins fang2024molinstructions SL 6.7B 0.01 0.705 31.23 0.283 0.487 0.23 1.00
HIGHT chen2024hight SL 6.7B 0.20 0.914 20.20 0.772 0.623 0.58 0.99
InstructMol cao2023instructmol SL 6.7B 0.41 0.941 13.97 0.753 0.852 0.71 1.00
PRESTO cao-etal-2024-presto GL 3.2B 0.53 0.958 10.30 0.823 0.887 0.79 1.00
Omni-Mol GL 2.2B 0.57 0.960 8.97 0.864 0.909 0.83 1.00
Model #Par Exa BLEU Lev RDK MAC Mor Val
@@
DeepSeekV3 zheng2023judging ICL 685B 0.26 0.684 15.20 0.501 0.581 0.47 0.99
Llama2 touvron2023llama SL 6.7B 0.00 0.283 53.51 0.136 0.294 0.11 1.00
Mol-Ins fang2024molinstructions SL 6.7B 0.04 0.224 23.17 0.237 0.364 0.21 1.00
HIGHT chen2024hight SL 6.7B 0.07 0.482 27.17 0.462 0.346 0.30 1.00
InstructMol cao2023instructmol SL 6.7B 0.13 0.610 19.66 0.444 0.539 0.40 1.00
PRESTO cao-etal-2024-presto GL 3.2B 0.21 0.712 16.31 0.544 0.607 0.48 1.00
Omni-Mol GL 2.2B 0.23 0.726 14.59 0.557 0.627 0.52 1.00
Model #Par Exa BLEU Lev RDK MAC Mor Val
Catalyst Prediction
DeepSeekV3 touvron2023llama ICL 685B 0.28 0.189 7.83 0.510 0.481 0.29 0.99
Vicuna v1.5 zheng2023judging SL 6.7B 0.69 0.703 2.45 0.883 0.869 0.69 1.00
nach0-base - - 0.00 0.07 36.44 0.129 0.055 0.01 0.85
Mol-Ins fang2024molinstructions SL 6.7B 0.00 0.110 28.42 0.031 0.045 0.02 0.99
T5Chem GL - 0.02 0.346 13.41 0.146 0.268 0.20 0.99
PRESTO cao-etal-2024-presto GL 6.7B 0.77 0.814 1.76 0.914 0.895 0.77 1.00
Omni-Mol GL 2.2B 0.72 0.792 1.96 0.904 0.886 0.72 1.00
Model Type #Param HOMO LUMO GAP Avg.
Quantum Mechanics Property Prediction Task
DeepSeekV3 ICL 685B 0.0200 0.0599 0.0457 0.0456
LLaMA2 ICL 6.7B 0.7367 0.8641 0.5152 0.7510
Vicuna ICL 13B 0.7135 3.6807 1.5407 1.9783
Mol-Ins SL 6.7B 0.0210 0.0210 0.0203 0.0210
HIGHT SL 6.7B 0.0056 0.0065 0.0077 0.0066
InstructMol SL 6.7B 0.0048 0.0050 0.0061 0.0050
Omni-Mol GL 2.2B 0.0038 0.0047 0.0049 0.0044
@@
DeepSeekV3 ICL 685B 0.181 0.095 0.319 0.133 0.249 0.231
GPT-4-0314 RT - 0.607 0.525 0.634 0.476 0.562 0.610
BioMedGPT GL 10B 0.234 0.141 0.386 0.206 0.332 0.308
Mol-Ins SL 6.7B 0.249 0.171 0.331 0.203 0.289 0.271
HIGHT SL 6.7B 0.498 0.397 0.582 0.414 0.518 0.525
InstructMol SL 6.7B 0.475 0.371 0.566 0.394 0.502 0.509
Omni-Mol GL 2.2B 0.529 0.440 0.604 0.447 0.541 0.571
Model #Par B-2 B-4 R-1 R-2 R-L M
@@
DeepSeekV3 ICL 685B 0.39 0.31 0.50 0.34 0.46 0.54
Llama2 SL 6.7B 0.28 0.23 0.35 0.22 0.30 0.47
3D-MoLM(S) SL 6.7B 0.32 0.26 0.40 0.26 0.35 0.52
3D-MoLM(G) GL 6.7B 0.32 0.26 0.40 0.26 0.35 0.52
Omni-Mol GL 2.2B 0.52 0.44 0.53 0.38 0.49 0.58
Model #Par B-H S-M
Yield Prediction
DeepSeekV3 ICL 685B -0.55 -1.09
Llama2 SL 6.7B -0.48 0.12
Vicuna v1.5 SL 6.7B -0.13 0.15
PRESTO GL 6.7B 0.94 0.65
Omni-Mol GL 2.2B 0.94 0.68
Model #Par Weight LogP TPSA
More Mol2Num
DeepSeekV3 ICL 685B 98.63(100) 2.26(100) 60.32(100)
Llama2 SL 6.7B 22.10 (96) 1.45 (95) 15.87 (92)
3D-MoLM(S) SL 6.7B 14.79 (95) 0.66 (97) 9.71 (93)
3D-MoLM(G) GL 6.7B 16.58 (92) 0.78 (95) 10.90 (90)
Omni-Mol GL 2.2B 11.07(100) 0.49(100) 5.89(100)
Model #Par Exa BLEU Lev RDK MAC Mor Val
Solvent Prediction
DeepSeekV3 ICL 685B 0.06 0.471 5.08 0.196 0.248 0.15 0.99
Vicuna v1.5 SL 6.7B 0.32 0.436 3.81 0.459 0.486 0.43 1.00
nach0-base - - 0.00 0.072 36.44 0.129 0.055 0.01 0.85
Mol-Instruction SL 6.7B 0.00 0.155 25.12 0.030 0.122 0.04 1.00
T5Chem GL - 0.08 0.311 16.22 0.458 0.424 0.40 0.99
PRESTO GL 6.7B 0.42 0.695 2.76 0.529 0.547 0.51 0.91
Omni-Mol GL 2.2B 0.52 0.759 2.71 0.671 0.673 0.64 1.00
Model #Par B-2 B-4 R-1 R-2 R-L
Experimental Procedure
DeepSeekV3 ICL 685B - - - - -
TextChemT5 GL 220M 0.541 0.406 0.615 0.403 0.564
MolT5-Large SL 780M 0.545 0.410 0.625 0.409 0.572
Galactica SL 1.3B 0.535 0.395 0.609 0.386 0.552
MolCA, Galac SL 1.3B 0.549 0.415 0.625 0.404 0.570
ReactXT, Galac SL 1.3B 0.574 0.440 0.644 0.427 0.589
Omni-Mol GL 2.2B 0.572 0.448 0.532 0.274 0.464
Model #Param $\text{QED}\geq 0.6$ $\text{drd2}\geq 0.5$
$\text{QED}\geq 0.6$
$\text{drd2}\geq 0.5$
$\Delta\geq 0.4$
$\text{QED}\geq 0.6$
$\Delta\geq 0.4$
$\text{drd2}\geq 0.5$
$\Delta\geq 0.4$
$\text{QED}\geq 0.6$
$\text{drd2}\geq 0.5$
Molecule Editing
Llama3.2-1B* SL 1.2B 0.8846 0.0507 0.0311 0.4441 0.0360 0.0271
DeepSeekV3 ICL 685B 0.8750 0.0000 0.0023 0.7500 0.0000 0.0017
Omni-Mol GL 2.2B 0.9612 0.0653 0.0412 0.7913 0.0560 0.0341
Table 3: Comprehensive comparisons on Mol2Mol, Mol2Text and Mol2Num tasks. Par: Parameters, Lev: Levenshtein, MAC: MACCS, Mor: Morgan, Val: Validity, Avg.: Average. B-2: BLEU-2, B-4: BLEU-4, R-1: ROUGE-1, R-2: ROUGE-2, R-L: ROUGE-L, M: METEOR, ICL: In-Context Learning, SL: Specialist, GL: Generalist, RT: Retrieval. * means our re-implementation. 3D-MoLM(S) is the specialist version and 3D-MoLM(G) is the generalist version.
We aim to address the following concerns: (1) Compared with existing baselines, can Omni-Mol achieve the best performances on the comprehensive omni-molecular datasets with 16 tasks? (2) Is Omni-Mol a scalable framework with the capacity and potential to solve complex molecular tasks? (3) Are all key components of Omni-Mol essential for omni-molecular task learning? We begin by describing the experimental setup and then answer all the questions in the subsequent sections.
6.1 Experimental Setup
Baselines. To ensure a fair comparison, we first choose representative LLM-based models such as InstructMol and HIGHT, and also report several previous baselines, including Mol-Instruction, Llama, and Vicuna zheng2023judging . We also conduct 5 shot in-context learning test on powerful open-source models like DeepSeekV3. For datasets with fewer models, we re-implement PRESTO as a baseline.